OpenStudy (anonymous):

Use the periodic properties of the trigonometric functions to find the exact value of the expression

OpenStudy (anonymous):

\[\cos \left(\begin{matrix}10\pi \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):

i need help on how to do it will fan and medal